# Nakai-Moishezon criteria and the toric Thomas-Yau conjecture

### v1

@@ 8.1.2 Restriction to a curve
Then, by (8.2), the phase inequality appearing in (3.3),
@@ 8.1.3 Sub-bundle
Then, by (8.2), the phase inequality appearing in (3.3),
@@ 9 General toric manifolds
for $k\gg 1$.
@@
$K^{(0)}_{k[\omega_{0}]}\left(\Theta^{-1}_{k[\omega_{0}]}(e^{-\operatorname{i}[\omega_{0}]}),\Theta^{-1}_{k[\omega_{0}]}(\operatorname{ch}(L))\right)$
$=\sum_{p\in\operatorname{Crit}(W_{k[\omega_{0}]})}\frac{\Theta^{-1}_{k[\omega_{0}]}(e^{-\operatorname{i}[\omega_{0}]})\Theta^{-1}_{k[\omega_{0}]}(\operatorname{ch}(L))}{\prod_{i}x^{2}_{i}\det\nabla^{2}W_{k[\omega_{0}]}}\big{|}_{p}\,\frac{\Omega_{k[\omega_{0}]}}{\Omega_{0}}\big{|}_{p}.$
@@
gives a linear function
$\Psi_{p}\in\operatorname{Hom}(\operatorname{GM}(W(k\omega_{0}))|_{z=0},\mathbb{C})$
which is then induced by integration along a complex cycle $\Gamma_{p}$, with
$[\Gamma_{p}]\in H_{n}(\mathcal{Y}_{q_{k}},\{\operatorname{Re}(W(k\omega_{0})/z)\gg 0\};\mathbb{Z})\otimes\mathbb{C},$
that is, we have, in the sense of [9], Section 6.2,
@@
where
$\Omega^{(k)}:=\frac{\Omega_{k[\omega_{0}]}}{\Omega_{0}}\Theta^{-1}_{k[\omega_{0}]}(e^{-\operatorname{i}[\omega_{0}]})\in\operatorname{GM}(W(k\omega_{0})).$
So, setting

### v2

@@ 8.1.2 Restriction to a curve
Then, by (8.2), under our genericity assumption on $[\omega_{0}]$, the phase inequality appearing in (3.3),
@@ 8.1.3 Sub-bundle
Then, by (8.2), under our genericity assumption on $[\omega_{0}]$, the phase inequality appearing in (3.3),
@@ 9 General toric manifolds
for $k\gg 1$, as long as $[\omega_{0}]$ is generic in the sense that it does not lie in the union of finitely many proper analytic subvarieties of $H^{1,1}(X,\mathbb{R})$ determined by $c_{1}(L)$ (the union ranging through all toric irreducible toric subvarieties $V$).
@@
$K^{(0)}_{k[\omega_{0}]}\left(\Theta^{-1}_{k[\omega_{0}]}(e^{-\operatorname{i}[\omega_{0}]}),\Theta^{-1}_{k[\omega_{0}]}(\operatorname{ch}(L))\right)=\sum_{p\in\operatorname{Crit}(W_{k[\omega_{0}]})}\frac{\Theta^{-1}_{k[\omega_{0}]}(e^{-\operatorname{i}[\omega_{0}]})\Theta^{-1}_{k[\omega_{0}]}(\operatorname{ch}(L))}{\prod_{i}x^{2}_{i}\det\nabla^{2}W_{k[\omega_{0}]}}\big{|}_{p}\,\frac{\Omega_{k[\omega_{0}]}}{\Omega_{0}}\big{|}_{p}.$
@@
gives a linear function $\Psi_{p}\in\operatorname{Hom}(\operatorname{GM}(W(k\omega_{0}))|_{z=0},\mathbb{C})$ which is then induced by integration along a complex cycle $\Gamma_{p}$, with $[\Gamma_{p}]\in H_{n}(\mathcal{Y}_{q_{k}},\{\operatorname{Re}(W(k\omega_{0})/z)\gg 0\};\mathbb{Z})\otimes\mathbb{C}$, that is, we have, in the sense of [9], Section 6.2,
@@
where $\Omega^{(k)}:=\frac{\Omega_{k[\omega_{0}]}}{\Omega_{0}}\Theta^{-1}_{k[\omega_{0}]}(e^{-\operatorname{i}[\omega_{0}]})\in\operatorname{GM}(W(k\omega_{0}))$. So, setting